# Fisher-Orthogonal Projected Natural Gradient Descent for Continual Learning

### v1

@@ Abstract
Continual learning aims to enable neural networks to acquire new knowledge on sequential tasks. However, the key challenge in such settings is to learn new tasks without catastrophically forgetting previously learned tasks. We propose the Fisher-Orthogonal Projected Natural Gradient Descent (FOPNG) optimizer, which enforces Fisher-orthogonal constraints on parameter updates to preserve old task performance while learning new tasks. Unlike existing methods that operate in Euclidean parameter space, FOPNG projects gradients onto the Fisher-orthogonal complement of previous task gradients. This approach unifies natural gradient descent with orthogonal gradient methods within an information-geometric framework. The resulting update direction is invariant under reparameterization, guarantees descent in the Fisher metric, and helps preserve prior task outputs. We provide theoretical analysis establishing the properties of the projected update, describe efficient and practical implementations using the diagonal Fisher, and demonstrate strong results on standard continual learning benchmarks such as Permuted-MNIST, Split-MNIST, Rotated-MNIST, Split-CIFAR10, and Split-CIFAR100.
@@ 2.6 Orthogonal Gradient Methods
While conceptually appealing, OGD inherits Euclidean geometry limitations: directions can be Euclidean-orthogonal yet cause significant output interference, or vice versa. Yadav et al. (2025) combined natural gradients with OGD by first computing a projection against past gradients in Euclidean space and pre-conditioning the resulting projection by the Fisher information. However, they found that performance decreased with this formulation. We show empirically that the projection must occur in Fisher space to respect the natural gradient geometry. A more recent work by Lu and Armour (2025) uses similar ideas about Fisher orthogonality, but derives a different optimizer for training with large batch sizes and does not discuss continual learning. Our work explicitly formulates continual learning in Fisher-geometric terms, deriving closed-form solutions that unify natural gradients with orthogonal projections.
@@ 3.1 Preliminaries and Notation
We clarify some notation that will be used throughout the rest of the work. Let $\theta\in\mathbb{R}^{p}$ denote model parameters and $p_{\theta}(y|x)$ a smooth conditional model. For a data distribution $q(x)$, define the Fisher information matrix:
$\mathbb{E}_{x\sim q}\mathbb{E}_{y\sim p_{\theta}(\cdot|x)}\nabla_{\theta}\log p_{\theta}(y|x)\nabla_{\theta}\log p_{\theta}(y|x)^{\top}$ (10)
For any $u,v\in\mathbb{R}^{p}$, define the Fisher inner product and norm:
$\langle u,v\rangle_{F}:=u^{\top}Fv,\quad\|u\|_{F}:=\sqrt{u^{\top}Fu}.$ (11)
Let $L_{\text{new}}(\theta)$ denote the new-task loss with gradient $g=\nabla_{\theta}L_{\text{new}}(\theta)$, and let $G=[g_{1},\ldots,g_{m}]\in\mathbb{R}^{p\times m}$ collect gradients on previous tasks.
@@
    $L_{\text{new}}$ is twice continuously differentiable in a neighborhood of $\theta$.
@@
    The Fisher matrix is positive definite.
@@
    G has full column-rank.
In Section 4.2 we show that assumption 2 can be relaxed by using a regularization term. Assumption 3 always holds in practice as the number of stored gradients is many orders of magnitude less than the number of model parameters.
@@ 3.2 Problem Formulation
FOPNG addresses continual learning through two simultaneous desiderata formulated in Fisher-Riemannian geometry:
Fisher-Orthogonal Preservation.
Let $G=[g_{1},\ldots,g_{m}]\in\mathbb{R}^{p\times m}$ collect $m$ gradients from previous tasks and let $F_{\text{old}}$ be the Fisher matrix computed on data from the previous task. We require that the new update $v$ satisfies
$G^{\top}F_{\text{old}}v=0.$ (12)
This ensures that $v$ is orthogonal to old task gradients under the Fisher metric, meaning the update avoids perturbing the output distributions on previous tasks.
Natural Gradient Step on New Task.
Let $g=\nabla_{\theta}L_{\text{new}}(\theta)$ be the gradient of the new task loss and $F_{\text{new}}$ be the Fisher matrix on data from the new task. We constrain the Fisher norm of the update: $v^{\top}F_{\text{new}}v\leq\epsilon^{2}$. This ensures that the change in output distribution on the new task is controlled, independent of parameterization. We found in our experiments that without this trust-region constraint, performance on old tasks collapses almost instantly after training on new ones.
@@ 3.3 Constrained Optimization Problem
Combining these desiderata, we formulate FOPNG as the following constrained optimization problem, where $g$ is the current batch gradient computed via backpropagation during training:
$\max_{v}\quad$ $g^{\top}v$ (13)
s.t. $v^{\top}F_{\text{new}}v\leq\epsilon^{2},$ (14)
$G^{\top}F_{\text{old}}v=0.$ (15)
The objective (13) seeks to maximize progress on the new task. Constraint (14) bounds the step size in the Fisher metric of the new task. Constraint (15) enforces Fisher-orthogonality to previous task gradients to reduce forgetting.
@@ 3.4 Deriving the FOPNG Update
Under Assumption 3.1, the solution to the optimization problem (13)–(15) is
$v^{*}=\epsilon\frac{F_{\text{new}}^{-\frac{1}{2}}PF_{\text{new}}^{-\frac{1}{2}}g}{\sqrt{g^{\top}F_{\text{new}}^{-\frac{1}{2}}P^{\top}PF_{\text{new}}^{-\frac{1}{2}}g}},$ (16)
where the projection matrix is
$P=I-M\left(M^{\top}M\right)^{-1}M^{\top}.$ (17)
@@
$M=F_{\text{new}}^{-\frac{1}{2}}F_{\text{old}}G$ (18)
@@
The matrix $P$ projects onto the Fisher-orthogonal complement of the subspace spanned by previous task gradients, and the resulting direction is then scaled to have the specified Fisher norm.
We also verify that this update makes progress on the task loss with the following result.
Theorem 3.3 (Descent Direction).
The update $v^{*}$ from Theorem 3.2 is a descent direction for the new task loss, i.e., $g^{\top}v^{*}\geq 0$.
Proof.
See Appendix A.4. ∎
3.5 Fisher Natural Gradient (FNG): The Unconstrained Baseline
To contextualize the effect of Fisher-orthogonality, we now derive the Fisher Natural Gradient (FNG), which corresponds to the same optimization problem as FOPNG but without the Fisher-orthogonality constraints. This serves as a principled baseline: it preserves the Fisher geometry but does not attempt to prevent interference with previously learned tasks.
@@ 3.5 Fisher Natural Gradient (FNG): The Unconstrained Baseline
Consider the constrained optimization problem
$\max_{v\in\mathbb{R}^{p}}\;g^{\top}v\qquad\text{s.t.}\qquad v^{\top}F_{\mathrm{new}}v\leq\epsilon^{2}.$
The optimal update direction is
$v_{\mathrm{FNG}}\;=\;\epsilon\,\frac{F_{\mathrm{new}}^{-1}g}{\sqrt{\,g^{\top}F_{\mathrm{new}}^{-1}g\,}}.$
@@ 3.6 FOPNG-PreFisher: Another Variant
FOPNG maintains a moving exponential average of the Fisher’s at each task as $F_{\text{old}}$. However, we can alternatively use a separate $F_{\text{old}}$ for every task. Instead of storing raw gradients $g_{i}$ in the gradient buffer, we precompute Fisher-weighted gradients $\tilde{g}_{i}=F_{i}g_{i}$, where $F_{i}$ is the Fisher matrix computed after training on task $i$. Instead, we store $\tilde{G}=[F_{1}G_{1},\ldots,F_{k-1}G_{k-1}]$ where $G_{i}$ is a matrix containing the gradients on task $i$.
The constrained optimization problem becomes:
$\max_{v}\quad$ $g^{\top}v\quad\textit{s.t.}\quad v^{\top}F_{\text{new}}v\leq\epsilon^{2},\;\;\tilde{G}^{\top}v=0$ (19)
where the orthogonality constraint is now Euclidean. Following similar derivations as in Section 3.4, the solution is:
$v^{*}=\epsilon\frac{F_{\text{new}}^{-\frac{1}{2}}PF_{\text{new}}^{-\frac{1}{2}}g}{\sqrt{g^{\top}F_{\text{new}}^{-\frac{1}{2}}P^{\top}PF_{\text{new}}^{-\frac{1}{2}}g}},$ (20)
where the projection matrix is
$P=I-M\left(M^{\top}M\right)^{-1}M^{\top}.$ (21)
where
$M=F_{\text{new}}^{-\frac{1}{2}}\tilde{G}$ (22)
This formulation eliminates the $F_{\text{old}}$ term entirely from the per-iteration update which has an additional benefit of reducing the number of Fisher-matrix-vector products. This makes FOPNG-PreFisher particularly attractive for longer task sequences.
@@ 4.1 Algorithm Description
Algorithms 1 and 2 present the complete FOPNG and FOPNG-PreFisher procedures respectively. The algorithm maintains a collection $G$ of gradients from previous tasks. For each new task, it computes both the new-task Fisher matrix $F_{\text{new}}$ and the old-task Fisher matrix $F_{\text{old}}$. The projection ensures Fisher-orthogonality while the normalization ensures bounded Fisher-norm steps. Note that the $\epsilon$ variable introduced in the theoretical formulation is simply folded into the learning rate $\eta.$
@@
8:    Compute $v$ according to Equation 20
@@ 5.1 Experimental Design
  • Split-CIFAR100 (Zenke et al., 2017) (10 tasks): Each task contains 10 image classes. We train on classes 1-10 first, then 11-20, and so on. This also uses a multi-head architecture.
@@
To establish a consistent, strong baseline on the first task (for which there is no previous task to ”forget”), we train using standard optimizers without continual learning constraints. For all methods, we use SGD on the first task with the same learning rate as the optimizer. All methods then switch to their respective continual learning optimizers starting from task 2. For Split-CIFAR10, all optimizers use SGD with a learning rate of 1e31e-3 on the first task to establish a consistent baseline on this more challenging dataset. For Split-CIFAR100, all optimizers use SGD with a learning rate of 1e21e-2 on the first task.
@@ 5.3 Ablation Studies
Figure 2: In general, larger batch sizes increase performance and reduce variance. Shaded regions show 95% confidence intervals across 5 independent runs.
@@ 6 Conclusion
We have presented a geometrically principled optimizer for continual learning that enforces Fisher-orthogonality to previous task gradients while taking minimal Fisher-norm steps. The resulting direction is invariant under reparameterization, guarantees descent in the Fisher metric, and reduces changes in prior task outputs after training on new tasks. This approach unifies orthogonal and natural gradient methods within the Fisher-Riemannian geometry, providing a principled solution to the stability-plasticity dilemma in continual learning. We view our work as a mix of theoretical and empirical contributions that future work may build upon to create more salient deep learning optimizers. Notably, a key takeaway of our work is demonstrating that orthogonal projections in Fisher-space are beneficial for reducing catastrophic forgetting.
@@ Appendix A Theoretical Derivations
Form the Lagrangian with multipliers $\mu\geq 0$ and $\nu\in\mathbb{R}^{m}$:
$\mathcal{L}(v,\mu,\nu)=-g^{\top}v+\frac{\mu}{2}(v^{\top}F_{\text{new}}v-\epsilon^{2})\\ +\nu^{\top}(G^{\top}F_{\text{old}}v).$ (30)
The first-order condition $\nabla_{v}\mathcal{L}=0$ gives:
$-g+\mu F_{\text{new}}v+F_{\text{old}}G\nu$ $=0$
$\implies v$ $=\frac{1}{\mu}\left(F_{\text{new}}^{-1}g-F_{\text{new}}^{-1}F_{\text{old}}G\nu\right).$ (31)
Apply the constraint $G^{\top}F_{\text{old}}v=0$:
$G^{\top}F_{\text{old}}F_{\text{new}}^{-1}g\\ =G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G\nu.$ (32)
Solve for $\nu$:
$\nu=\left(G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G\right)^{-1}\\ \times G^{\top}F_{\text{old}}F_{\text{new}}^{-1}g.$ (33)
Let $M=F_{\text{new}}^{-\frac{1}{2}}F_{\text{old}}G$. Define
$P=I-M(M^{\top}M)M^{\top}.$ (34)
Then:
$v=\frac{1}{\mu}F_{\text{new}}^{-\frac{1}{2}}PF_{\text{new}}^{-\frac{1}{2}}g.$ (35)
Enforce $v^{\top}F_{\text{new}}v=\epsilon^{2}$:
$\frac{1}{\mu^{2}}g^{\top}F_{\text{new}}^{-\frac{1}{2}}P^{\top}F_{\text{new}}^{-\frac{1}{2}}F_{\text{new}}F_{\text{new}}^{-\frac{1}{2}}PF_{\text{new}}^{-\frac{1}{2}}g$ $=\epsilon^{2}$
$\implies\mu=\frac{\sqrt{g^{\top}F_{\text{new}}^{-\frac{1}{2}}P^{\top}PF_{\text{new}}^{-\frac{1}{2}}g}}{\epsilon}.$ (36)
The solution is:
$v^{*}=\epsilon\frac{F_{\text{new}}^{-\frac{1}{2}}PF_{\text{new}}^{-\frac{1}{2}}g}{\sqrt{g^{\top}F_{\text{new}}^{-\frac{1}{2}}P^{\top}PF_{\text{new}}^{-\frac{1}{2}}g}}.$ (37)
A.4 Proof of Theorem 3.3: Descent Direction
@@
We have $v^{*}=cF_{\text{new}}^{-\frac{1}{2}}PF_{\text{new}}^{-\frac{1}{2}}g$ for positive scalar $c=\epsilon/\sqrt{g^{\top}P^{\top}F_{\text{new}}^{-1}Pg}$. Under the transformation $w=F_{\text{new}}^{-\frac{1}{2}}g$ this becomes
$g^{\top}v^{*}=c(w^{\top}w-w^{\top}M\left(M^{\top}M\right)^{-1}M^{\top}w)$ (38)
We can recognize the second term as a projection of $w$ under matrix $M$, hence $w^{\top}w\geq w^{\top}M\left(M^{\top}M\right)M^{\top}w.$ Therefore we have $g^{\top}v^{*}\geq 0$. ∎
A.5 Proof of Theorem 3.4: Fisher Natural Gradient
Proof.
Consider the Lagrangian
$\mathcal{L}(v,\lambda)=g^{\top}v\\ -\mu\left(v^{\top}F_{\mathrm{new}}v-\epsilon^{2}\right),\qquad\lambda\geq 0.$ (39)
Taking the derivative with respect to $v$ and setting it to zero gives
$\nabla_{v}\mathcal{L}$ $=g-2\mu F_{\mathrm{new}}v=0$
$\implies v=\frac{1}{2\mu}F_{\mathrm{new}}^{-1}g.$ (40)
At the optimum the constraint is active, so
$v^{\top}F_{\mathrm{new}}v$ $=\left(\frac{1}{2\mu}\right)^{2}g^{\top}F_{\mathrm{new}}^{-1}g=\epsilon^{2}.$ (41)
Solving for $(2\mu)^{-1}$ yields
$\frac{1}{2\mu}=\frac{\epsilon}{\sqrt{g^{\top}F_{\mathrm{new}}^{-1}g}}.$ (42)
Substituting this expression back into $v$ gives the optimal solution
$v_{\mathrm{FNG}}=\epsilon\,\frac{F_{\mathrm{new}}^{-1}g}{\sqrt{g^{\top}F_{\mathrm{new}}^{-1}g}}.$ (43)

### v2

@@ Abstract
Continual learning aims to enable neural networks to acquire new knowledge on sequential tasks. However, the key challenge in such settings is to learn new tasks without catastrophically forgetting previously learned tasks. We propose the Fisher-Orthogonal Projected Natural Gradient Descent (FOPNG) optimizer, which enforces Fisher-orthogonal constraints on parameter updates to preserve old task performance while learning new tasks. Unlike existing methods that operate in Euclidean parameter space, FOPNG projects gradients onto the Fisher-orthogonal complement of previous task gradients. This approach unifies natural gradient descent with orthogonal gradient methods within an information-geometric framework. We provide theoretical analysis deriving the projected update, describe efficient and practical implementations using the diagonal Fisher, and demonstrate strong results on standard continual learning benchmarks such as Permuted-MNIST, Split-MNIST, Rotated-MNIST, Split-CIFAR10, and Split-CIFAR100. Our code is available at https://github.com/ishirgarg/FOPNG.
@@ 2.6 Orthogonal Gradient Methods
While conceptually appealing, OGD inherits Euclidean geometry limitations: normal gradients do not capture the effect on the output distribution of a model. Yadav et al. (2025) combined natural gradients with OGD by first computing a projection against past gradients in Euclidean space and pre-conditioning the resulting projection by the Fisher information. However, they found that performance decreased with this formulation. We show empirically that the projection must occur in Fisher space to respect the natural gradient geometry. A more recent work by Lu and Armour (2025) uses similar ideas about Fisher orthogonality, but derives a different optimizer for training with large batch sizes and does not discuss continual learning. Our work explicitly formulates continual learning in Fisher-geometric terms, deriving closed-form solutions that unify natural gradients with orthogonal projections.
@@ 3.1 Preliminaries and Notation
Let $\theta\in\mathbb{R}^{p}$ denote model parameters and $p_{\theta}(y\mid x)$ a smooth conditional distribution. For a data distribution $q(x)$, define the Fisher information matrix
$F=\mathbb{E}_{x\sim q}\mathbb{E}_{y\sim p_{\theta}(\cdot\mid x)}\big[\nabla_{\theta}\log p_{\theta}(y\mid x)\nabla_{\theta}\log p_{\theta}(y\mid x)^{\top}\big].$ (10)
For any $u,v\in\mathbb{R}^{p}$, define the Fisher inner product and norm
$\langle u,v\rangle_{F}:=u^{\top}Fv,\qquad\|u\|_{F}:=\sqrt{u^{\top}Fu}.$ (11)
Let $L_{\text{new}}(\theta)$ be the new-task loss and $g=\nabla_{\theta}L_{\text{new}}(\theta)$ its gradient. Let $G=[g_{1},\ldots,g_{m}]\in\mathbb{R}^{p\times m}$ collect gradients from previous tasks. We denote by $F_{\text{new}}$ and $F_{\text{old}}$ the Fisher matrices computed on new-task and previous-task data, respectively.
@@
    The score function $\nabla_{\theta}\log p_{\theta}(y\mid x)$ exists and has finite second moments, so that the Fisher information matrices are well-defined.
@@
    $F_{\text{new}}$ is positive definite.
@@
    $G$ has full column rank.
In Section 4.2 we show that Assumption 2 can be relaxed using standard regularization techniques. Assumption 3 holds in practice since the number of stored gradients is far smaller than the parameter dimension.
3.2 Geometric Motivation
In information geometry, the natural gradient $F_{\text{new}}^{-1}g$ represents the steepest descent direction for the new-task loss under the Fisher metric. Any learning rule that aims to be invariant to reparameterization should therefore be expressed relative to this geometry.
We adopt the viewpoint that learning proceeds by selecting a Fisher-preconditioned direction that approximates the raw gradient $g$ while respecting constraints imposed by previously learned tasks. To formalize this idea, we introduce an auxiliary variable $u\in\mathbb{R}^{p}$ and define the actual parameter update for some scalar $c>0$ as
$v=cF_{\text{new}}^{-1}(g-u).$ (12)
The vector $u$ represents the component of the gradient that must be removed or modified in order to satisfy constraints derived from prior tasks.
Crucially, we reason about $u$ through its Fisher-preconditioned form $F_{\text{new}}^{-1}u$, which lives in the same geometric space as the natural gradient.
@@ 3.3 Problem Formulation
FOPNG is defined by three principles, each motivated by Fisher geometry.
Natural-Gradient Approximation.
We seek a Fisher-preconditioned direction $F_{\text{new}}^{-1}u$ that approximates the raw gradient $g$ as closely as possible. This is captured by minimizing the Fisher norm of the discrepancy:
$\min_{u}\|g-F_{\text{new}}^{-1}u\|_{F_{\text{new}}}^{2}.$ (13)
This objective measures how well the chosen direction aligns with the natural gradient of the new task.
Compatibility with Previous Tasks.
To limit interference with previously learned tasks, we require that the Fisher-conditioned component of $u$ aligns with directions spanned by stored gradients:
$F_{\text{old}}^{-1}u\in\operatorname{Col}(G).$ (14)
Intuitively, we can think about $u$ as telling us how much of the previous task gradients are contained in the natural gradient update for $u$. FOPNG attempts to remove the old-task interfering component from the new task gradient.
Natural-Gradient Trust Region.
Finally, we constrain the magnitude of the update in the Fisher metric of the new task. For some learning rate $\eta>0,$ we enforce
$v||_{F_{\text{new}}}=\eta.$ (15)
This trust-region constraint bounds the KL divergence induced by the natural gradient update and helps control the learning rate across tasks. Empirically, we find that this is needed for learning stability.
@@ 3.4 Constrained Optimization Problem
Combining these principles yields the following constrained optimization problem:
$\min_{u\in\mathbb{R}^{p}}\quad$ $\|g-F_{\text{new}}^{-1}u\|_{F_{\text{new}}}^{2}$ (16)
s.t. $F_{\text{old}}^{-1}u\in\operatorname{Col}(G),$ (17)
The resulting update is $v=cF_{\text{new}}^{-1}(g-u)$ where $c$ is chosen to satisfy Equation 15.
@@ 3.5 Deriving the FOPNG Update
Under Assumption 3.1, the FOPNG update is
$v^{*}=\eta\frac{F_{\text{new}}^{-1}Pg}{\sqrt{g^{\top}P^{\top}F_{\text{new}}^{-1}Pg}},$ (18)
@@
$P=I-F_{\text{old}}G(G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G)^{-1}G^{\top}F_{\text{old}},$ (19)
@@
The matrix $P$ projects onto the subspace that is Fisher-orthogonal to previous-task gradients when expressed in the geometry of the new task. The final scaling enforces the natural-gradient trust region.
3.6 Fisher Natural Gradient (FNG): An Unconstrained Baseline
In order to evaluate the effect of our projections, we derive a simpler natural-gradient update that removes the projection constraint.
If the compatibility constraint $F_{\text{old}}^{-1}u\in\operatorname{Col}(G)$ is removed, the problem reduces to selecting the best Fisher-conditioned approximation to $g$ under the trust-region constraint. This effectively yields the standard natural gradient descent (Amari, 1998).
@@ 3.6 Fisher Natural Gradient (FNG): An Unconstrained Baseline
Consider the constraints
$v_{\mathrm{FNG}}$ $=cF_{\mathrm{new}}^{-1}g$ (20)
[…]
The resulting update is
$v_{\mathrm{FNG}}=\eta\frac{F_{\mathrm{new}}^{-1}g}{\sqrt{g^{\top}F_{\mathrm{new}}^{-1}g}}$
@@ 3.7 FOPNG-PreFisher: Another Variant
Instead of maintaining a single aggregate Fisher matrix $F_{\text{old}}$, we may store task-specific Fisher-weighted gradients. For each previous task $i$, define $\tilde{g}_{i}=F_{i}g_{i}$ and collect them as
$\tilde{G}=[F_{1}G_{1},\ldots,F_{k-1}G_{k-1}].$
The optimization problem becomes
$\min_{u\in\mathbb{R}^{p}}\quad$ $\|g-F_{\text{new}}^{-1}u\|_{F_{\text{new}}}^{2}$ (22)
[…]
The solution has the same form as (18), with
$P=I-\tilde{G}(\tilde{G}^{\top}F_{\text{new}}^{-1}\tilde{G})^{-1}\tilde{G}^{\top}$ (24)
This variant removes $F_{\text{old}}$ from the per-iteration update and reduces the number of Fisher matrix–vector products, making it particularly attractive for long task sequences. We call this variation the FOPNG-PreFisher update.
@@ 4.1 Algorithm Description
Algorithms 1 and 2 present the complete FOPNG and FOPNG-PreFisher procedures respectively. The algorithm maintains a collection $G$ of gradients from previous tasks. For each new task, it computes both the new-task Fisher matrix $F_{\text{new}}$ and the old-task Fisher matrix $F_{\text{old}}$. The projection ensures Fisher-orthogonality while the normalization ensures bounded Fisher-norm steps.
@@
8:    Compute $v$ according to Equations 18, 24
@@ 5.1 Experimental Design
  • Split-CIFAR100 (Zenke et al., 2017) (10 tasks): Each task contains 10 image classes out of 100 total. We train on classes 1-10 first, then 11-20, and so on. This also uses a multi-head architecture.
@@
To establish a consistent, strong baseline on the first task (for which there is no previous task to ”forget”), we train using standard optimizers without continual learning constraints. For MNIST datasets, we use SGD on the first task with the same learning rate as the optimizer; due to the simplicity of MNIST this is able to learn the first task with almost perfect accuracy. All methods then switch to their respective continual learning optimizers starting from task 2. For Split-CIFAR10, all optimizers use SGD with a learning rate of 1e31e-3 on the first task to establish a consistent baseline on this more challenging dataset. For Split-CIFAR100, all optimizers use SGD with a learning rate of 1e21e-2 on the first task.
@@ 5.3 Ablation Studies
Figure 2: Larger batch sizes increase performance and reduce variance. Shaded regions show 95% confidence intervals across 5 independent runs.
@@ 6 Conclusion
We have presented a geometrically principled optimizer for continual learning that enforces Fisher-orthogonality to previous task gradients while taking minimal Fisher-norm steps. The resulting direction is invariant under reparameterization, guarantees descent in the Fisher metric, and reduces changes in prior task outputs after training on new tasks. This approach unifies orthogonal and natural gradient methods within the Fisher-Riemannian geometry, providing a principled solution to the stability-plasticity dilemma in continual learning. We view our work as a mix of theoretical and empirical contributions that future work may build upon to create more salient deep learning optimizers. Notably, a key takeaway of our work is demonstrating that projections in Fisher-space are beneficial for reducing catastrophic forgetting.
@@ Appendix A Theoretical Derivations
We solve the constrained optimization problem
$\min_{u\in\mathbb{R}^{p}}\quad$ $\|g-F_{\text{new}}^{-1}u\|_{F_{\text{new}}}^{2}$ (32)
s.t. $F_{\text{old}}^{-1}u\in\operatorname{Col}(G).$ (33)
Expand the objective:
$\|g-F_{\text{new}}^{-1}u\|_{F_{\text{new}}}^{2}$ $=(g-F_{\text{new}}^{-1}u)^{\top}F_{\text{new}}(g-F_{\text{new}}^{-1}u)$
$=g^{\top}F_{\text{new}}g-2g^{\top}u+u^{\top}F_{\text{new}}^{-1}u.$ (34)
The constraint $F_{\text{old}}^{-1}u\in\operatorname{Col}(G)$ means there exists $\alpha\in\mathbb{R}^{m}$ such that
$F_{\text{old}}^{-1}u=G\alpha\quad\Longleftrightarrow\quad u=F_{\text{old}}G\alpha.$ (35)
Substitute into the objective and minimize over $\alpha$:
$\min_{\alpha}\quad g^{\top}F_{\text{new}}g-2g^{\top}F_{\text{old}}G\alpha\\ +\alpha^{\top}G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G\alpha.$ (36)
Taking the derivative with respect to $\alpha$ and setting to zero:
$-2G^{\top}F_{\text{old}}g+2G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G\alpha$ $=0$
$\implies\alpha^{*}=(G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G)^{-1}G^{\top}F_{\text{old}}g.$ (37)
The optimal $u^{*}$ is therefore
$u^{*}=F_{\text{old}}G(G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G)^{-1}G^{\top}F_{\text{old}}g.$ (38)
Define the projection matrix
$P=I-F_{\text{old}}G(G^{\top}F_{\text{old}}F_{\text{new}}^{-1}F_{\text{old}}G)^{-1}G^{\top}F_{\text{old}}.$ (39)
Then $g-u^{*}=Pg$, and the unnormalized update is
$\tilde{v}=cF_{\text{new}}^{-1}(g-u^{*})=cF_{\text{new}}^{-1}Pg.$ (40)
To enforce the trust region constraint $v^{\top}F_{\text{new}}v=\eta^{2}$, we compute
$\tilde{v}^{\top}F_{\text{new}}\tilde{v}$ $=c^{2}(Pg)^{\top}F_{\text{new}}^{-1}F_{\text{new}}F_{\text{new}}^{-1}(Pg)$
$=c^{2}g^{\top}P^{\top}F_{\text{new}}^{-1}Pg.$ (41)
Setting this equal to $\eta^{2}$ gives
$c=\frac{\eta}{\sqrt{g^{\top}P^{\top}F_{\text{new}}^{-1}Pg}}.$ (42)
The final update is
$v^{*}=\eta\frac{F_{\text{new}}^{-1}Pg}{\sqrt{g^{\top}P^{\top}F_{\text{new}}^{-1}Pg}}.$ (43)
A.4 Proof of Theorem 3.3: Fisher Natural Gradient
@@
We parameterize the update as $v=cF_{\mathrm{new}}^{-1}g$ and seek the value of $c$ satisfying the trust region constraint $v^{\top}F_{\mathrm{new}}v=\eta^{2}$.
Substituting the parameterization:
$v^{\top}F_{\mathrm{new}}v$ $=(cF_{\mathrm{new}}^{-1}g)^{\top}F_{\mathrm{new}}(cF_{\mathrm{new}}^{-1}g)$
$=c^{2}g^{\top}F_{\mathrm{new}}^{-1}g.$ (44)
Setting this equal to $\eta^{2}$:
$c^{2}g^{\top}F_{\mathrm{new}}^{-1}g=\eta^{2}.$ (45)
Solving for $c$:
$c=\frac{\eta}{\sqrt{g^{\top}F_{\mathrm{new}}^{-1}g}}.$ (46)
Substituting back into the parameterization gives the FNG update:
$v_{\mathrm{FNG}}=\eta\,\frac{F_{\mathrm{new}}^{-1}g}{\sqrt{g^{\top}F_{\mathrm{new}}^{-1}g}}.$ (47)